# Bending parameterization of one-sided degenerated Kleinian surface groups

### v1

@@ 4.5. Geometric convergence
Let $(\rho_{n})_{n}$ be a sequence of representations in ${\mathcal{AH}}_{o}^{+}(S;e^{-})$ as in Proposition 4.3 with $\operatorname{m}^{-}\times\operatorname{b}^{+}(\rho_{n})=(m^{-}_{n},\lambda^{+}_{n})$ converging to $(m^{-},\lambda^{+})$ in ${\mathcal{T}}(\Sigma^{-})\times{\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$. By Proposition 4.3, up to replacing $(\rho_{n})_{n}$ by a subsequence, we assume that it converges algebraically to a representation $\rho_{\infty}\in{\mathcal{AH}}_{o}(S)$. The goal of this section is to show that $\rho_{\infty}$ belongs to ${\mathcal{AH}}_{o}^{+}(S;e^{-})$, i.e. that ${\mathbb{H}}^{3}/\rho_{\infty}(\Gamma)$ has parabolic locus $P$, no new geometrically infinite end, and that it still has $L$ as its ending lamination.
@@
The top end can be controlled thanks to the following result of Bonahon-Otal [BO04, §5,6] and Lecuire [Lec04, §4.1], [Lec06, §4.2]. We state a slightly weaker version, although their result also works in presence of parabolics in the top end.
@@
Let $(f_{n},\pi_{1}(S),\rho_{n})_{n}$ be a sequence of convex pleated surfaces such that $(\rho_{n})_{n}$ converges algebraically to some $\rho_{\infty}$. Let $(\lambda_{n})_{n}$ be the sequence of bending laminations of those pleated surfaces. Assume that $\lambda_{n}\to\lambda$ in ${\mathcal{ML}}(S)$ and that the closed leaves of $\lambda_{n}$ and $\lambda$ have weight $<\pi$. Then $(f_{n},\pi_{1}(S),\rho_{n})_{n}$ converges geometrically to a convex pleated surface $(f_{\infty},\pi_{1}(S),\rho_{\infty})$ inside ${\mathbb{H}}^{3}/\rho_{\infty}(\pi_{1}(S))$. In particular, the top boundary of the convex core of ${\mathbb{H}}^{3}/\rho_{\infty}(\pi_{1}(S))$ is homeomorphic to $S^{+}$.
@@ 5. Real-analyticity
Pick a maximal geodesic lamination $L$ containing the support of $\lambda^{+}$. By Bonahon’s shear-bend parameterization Theorem 2.5, the open subset ${\mathcal{R}}(L)\subset{\mathcal{X}}(S)$ consisting of those representations realizing a pleated surface with pleating locus contained in $\lambda^{+}$ is biholomorphically homeomorphic to an open subset of ${\mathcal{H}}(L;{\mathbb{C}}/2\pi i{\mathbb{Z}})$ via the shear-bend coordinates $\sigma+i\beta$. For a generic $\rho\in{\mathcal{R}}(L)$, the associated pleated surface $P_{\rho}$ lies inside the convex core of ${\mathbb{H}}^{3}/\rho(\pi_{1}(S))$ and is neither convex nor concave, in particular $\beta(\rho)$ takes positive and negative values in $[-\pi,\pi]$, i.e. it does not correspond to a transverse measure. However, if $\rho\in{\mathcal{R}}(L)$ does realize $P_{\rho}$ as the upper boundary of the convex core of ${\mathbb{H}}^{3}/\rho(\pi_{1}(S))$, then $\beta(\rho)$ corresponds to a transverse measure with support contained in $L$. In particular, for $\rho_{0}\in{\mathcal{AH}}_{o}^{+}(S;e^{-})$ with $\operatorname{b}^{+}(\rho_{0})=\lambda^{+}$, the real-analytic subset $\beta^{-1}(\beta(\rho_{0}))\subset{\mathcal{R}}(L)$ consists of those representations realizing a pleated surface with the same bending cocycle as $\rho_{0}$, that is, in this case, a convex pleated surface with bending lamination $\lambda^{+}$. In other words, $(\operatorname{b}^{+})^{-1}(\lambda^{+})\cap{\mathcal{R}}(L)=\beta^{-1}(\beta(\rho_{0}))\cap{\mathcal{R}}(L)$. Since ${\mathcal{R}}(L)$ is open in ${\mathcal{X}}(S)$, this shows that $(\operatorname{b}^{+})^{-1}(\lambda^{+})$ is locally real-analytic around $\rho_{0}$. Thus it is real-analytic. ∎
@@ 6.2. Prescribed third fundamental form
Given $\rho\in{\mathcal{QF}}_{o}(S)$, a theorem of Labourie [Lab92] states that each end of ${\mathbb{H}}^{3}/\rho(\pi_{1}(S))$ is uniquely foliated by smooth closed subsurfaces of constant curvature equal to $K$, with $K$ contained in the interval $(-1,0)$. The third fundamental form of such a subsurface has constant curvature $K^{*}\coloneqq K/(K+1)$. Thus it gives a point in the rescaled Teichmüller space ${\mathcal{T}}^{K^{*}}(S)$, which is the space of isotopy classes of metrics of constant curvature $K^{*}$ on $S$. Focusing on the top end, it follows that, for each $K\in(-1,0)$, one can define the map
(6) $III^{+}_{K}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{T}}^{K^{*}}(S)$
sending a quasifuchsian representation $\rho$ to the third fundamental form of the constant curvature $K$ surface in the top end of ${\mathbb{H}}^{3}/\rho(\pi_{1}(S))$. The third fundamental form $III^{+}_{K}$ is a “smooth measure of bending” and Chen-Schlenker proved that the analogous “smooth version” of Proposition 6.1 holds (see Corollary 1.4 and the proof of Lemma 4.1 in [CS25]). We need the following special case of their result, which is more general.
@@
The map $\operatorname{m}^{-}\times III^{+}_{K}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{T}}(S^{-})\times{\mathcal{T}}^{K^{*}}(S^{+})$ is a homeomorphism.
This is a mixed version of the earlier result of Schlenker that $III^{-}_{K}\times III^{+}_{K}$ is a homeomorphism [Sch06], which was an essential ingredient in [DS24].
@@
When $K\to-1$, the maps $III^{+}_{K}$ converge to $\operatorname{b}^{+}$ in the length spectrum, uniformly on compacts. More precisely, $\ell_{III^{+}_{K}(\rho)}(c)\to i(\lambda^{+}(\rho),c)$ when $K\to-1$, for each closed curve $c$ in $S$ and each $\rho\in{\mathcal{QF}}_{o}(S)$, and the convergence is uniform for $\rho$ in a compact subset of ${\mathcal{QF}}_{o}(S)$.
In order to apply the variant of Finney’s theorem as in the next section, we identify the targets of the maps $III^{+}_{K}$, for $K\in(-1,0)$, and $\operatorname{b}^{+}$, via the earthquake map $E\colon{\mathcal{ML}}(S)\times{\mathcal{T}}(S)\to{\mathcal{T}}(S)\colon(\lambda,m)\mapsto E(\lambda,m)$ as in [DS24, §4.1]. More precisely, we fix a point $m_{0}\in{\mathcal{T}}(S)$ and consider, for each $K\in(-1,0)$, the map
(7) $V_{K}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{ML}}(S^{+})\colon\rho\mapsto\frac{1}{\sqrt{\left|K^{*}\right|}}E(-,m_{0})^{-1}(\left|K^{*}\right|III^{+}_{K}(\rho)),$
@@ 6.6. Fibres of a limit of homeomorphism
Let $\operatorname{\mathbb{S}}^{k}\subset{\mathbb{D}}^{k+1}$ denote the inclusion of the $k$-sphere as the boundary of the $(k+1)$-ball. Recall that a space $K$ is weakly contractible if, for any $k\geq 0$, any map $\operatorname{\mathbb{S}}^{k}\to K$ extends to a map ${\mathbb{D}}^{k+1}\to K$. If a space $X$ is weakly contractible, then it is acyclic, i.e. all its reduced homology groups are trivial.
@@
Let $N$ and $N^{\prime}$ be two manifolds of the same dimension. Let $(f_{n}\colon N\to N^{\prime})_{n\in{\mathbb{N}}}$ be a sequence of open embeddings converging continuously (equivalently, compactly) to the continuous map $f_{\infty}\colon N\to N^{\prime}$, i.e. $f_{n}(x_{n})\to f_{\infty}(x_{\infty})$ whenever $x_{n}\to x_{\infty}$. If a fibre of $f_{\infty}$ is compact and is an absolute neighbourhood retract, then it is weakly contractible.
In fact, a compact weakly contractible ANR is automatically contractible, although we only need the fact that it is acyclic.
@@
Let $K\coloneqq f_{\infty}^{-1}(\ast)$ be a compact fibre, which is a retract of an open neighbourhood $\Omega$, which we can assume to have compact closure. Assuming that $K$ is not weakly contractible, we will show that $f_{\infty}$ sends a point of $\partial\Omega$ to $\ast$.
By contradiction, assume that there exists $k\geq 0$ and a map $\sigma\colon S^{k}\to K$, which is not contractible, i.e. does not extend to a map ${\mathbb{D}}^{k+1}\to K$. In particular, there is also no map ${\mathbb{D}}^{k+1}\to\Omega$ extending $\sigma$, for one could compose it with a retraction $\Omega\to K$ to obtain a map which lands in $K$.
Let $U\subset N^{\prime}$ be a contractible neighbourhood of $\ast$. Since $K$ is compact, we have $f_{n}(K)\subset U$ for all $n$ large enough. Fix such a $n\in{\mathbb{N}}$. Thus, $f_{n}(K)\subset U$. Since $U$ is contractible, the map $f_{n}\circ\sigma\colon\operatorname{\mathbb{S}}^{k}\to U$ extends to a map $g\colon{\mathbb{D}}^{k+1}\to U$. But recall that $f_{n}$ is an open embedding, hence $f_{n}\circ\sigma$ is not contractible inside $f_{n}(\Omega)$. Thus $g({\mathbb{D}}^{k+1})$ cannot be entirely contained in $f_{n}(\Omega)$, so it must intersect $f_{n}(\partial\Omega)$, say at a point $f_{n}(x_{U})$, for otherwise $g({\mathbb{D}}^{k+1})$ would be disconnected. In particular, $x_{U}\in\partial\Omega$ and $f_{n}(x_{U})\in U$.
@@
which is impossible since $\partial\Omega$ does not meet $K=f_{\infty}^{-1}(\ast)$. This concludes the proof that $K$ is weakly contractible. ∎
@@
Since $\operatorname{m}^{-}\times\operatorname{b}^{+}$ is the limit of the homeomorphisms $\operatorname{m}^{-}\times V_{K}$ and the convergence is uniform on compacts, $F$ is weakly contractible by Theorem 6.7, hence it is acyclic.
Since $F$ is a compact real-analytic space, it has a fundamental class modulo two (by Borel-Haefliger and Sullivan, see Theorem 5.3), i.e. a nonzero homology class $\alpha\in H_{\dim(F)}(F;{\mathbb{Z}}/2)$. But $F$ is acyclic (in fact, contractible), hence its homology is trivial in all degrees except 00. Since $\alpha\neq 0$, we must have $\dim(F)=0$. Therefore, $F$ is a contractible compact real-analytic space of dimension 00, that is, a point. This proves that $\operatorname{m}^{-}\times\operatorname{b}^{+}$ is injective. By invariance of domains, it is an open embedding. This concludes, since a proper open embedding into a connected space is a homeomorphism. ∎
@@ 6.8. Conclusion
is proper (Proposition 4.1), continuous (on the first factor, by the simultaneous uniformization theorem, see Section 2.10 ; on the second factor, by Proposition 3.4), and its domain and target are manifolds of the same dimension (By [Ker92, Theorem 4.7], <${\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ is a manifold of the same dimension as ${\mathcal{T}}(S^{+})$), it suffices to show that it is injective. With the notation of Section 1.1, it suffices to show that $\operatorname{b}^{+}_{\infty}$ is injective. But it is the limit of the open embeddings $\operatorname{b}^{+}_{t}$, when $t\to\infty$, and the convergence is continuous since $\operatorname{B}^{+}$ is continuous. By following the proof of Proposition 6.1 and using Theorem 6.7, we obtain that $\operatorname{b}^{+}_{\infty}$ is injective, hence $\operatorname{b}^{+}$ is injective too. ∎

### v2

@@ 4.5. Geometric convergence
[…]
@@
The top end can be controlled thanks to the following result of Bonahon-Otal [BO04, §5,6] and Lecuire [Lec25, §4.1], [Lec06, §4.2]. We state a slightly weaker version, although their result also works in presence of parabolics in the top end.
@@
Let $(f_{n},\pi_{1}(S),\rho_{n})_{n}$ be a sequence of convex pleated surfaces such that $(\rho_{n})_{n}$ converges algebraically to some $\rho_{\infty}$. Let $(\lambda_{n})_{n}$ be the sequence of bending laminations of those pleated surfaces. Assume that $\lambda_{n}\to\lambda$ in ${\mathcal{ML}}(S)$ and that the closed leaves of $\lambda_{n}$ and $\lambda$ have weight $\pi$. Then $(f_{n},\pi_{1}(S),\rho_{n})_{n}$ converges geometrically to a convex pleated surface $(f_{\infty},\pi_{1}(S),\rho_{\infty})$ inside ${\mathbb{H}}^{3}/\rho_{\infty}(\pi_{1}(S))$. In particular, the top boundary of the convex core of ${\mathbb{H}}^{3}/\rho_{\infty}(\pi_{1}(S))$ is homeomorphic to $S^{+}$.
@@ 5. Real-analyticity
Pick a maximal geodesic lamination $L$ containing the support of $\lambda^{+}$. By Bonahon’s shear-bend parameterization Theorem 2.5, the open subset ${\mathcal{R}}(L)\subset{\mathcal{X}}(S)$ consisting of those representations realizing a pleated surface with pleating locus contained in $\lambda^{+}$ is biholomorphically homeomorphic to an open subset of ${\mathcal{H}}(L;{\mathbb{C}}/2\pi i{\mathbb{Z}})$ via the shear-bend coordinates $\sigma+i\beta$. For a generic $\rho\in{\mathcal{R}}(L)$, the associated pleated surface $P_{\rho}$ lies inside the convex core of ${\mathbb{H}}^{3}/\rho(\pi_{1}(S))$ and is neither convex nor concave, in particular $\beta(\rho)$ takes positive and negative values in $\pi,\pi]$, i.e. it does not correspond to a transverse measure. However, if $\rho\in{\mathcal{R}}(L)$ does realize $P_{\rho}$ as the upper boundary of the convex core of ${\mathbb{H}}^{3}/\rho(\pi_{1}(S))$, then $\beta(\rho)$ corresponds to a transverse measure with support contained in $L$. In particular, for $\rho_{0}\in{\mathcal{AH}}_{o}^{+}(S;e^{-})$ with $\operatorname{b}^{+}(\rho_{0})=\lambda^{+}$, the real-analytic subset $\beta^{-1}(\beta(\rho_{0}))\subset{\mathcal{R}}(L)$ consists of those representations realizing a pleated surface with the same bending cocycle as $\rho_{0}$, that is, in this case, a convex pleated surface with bending lamination $\lambda^{+}$. In other words, $(\operatorname{b}^{+})^{-1}(\lambda^{+})\cap{\mathcal{R}}(L)=\beta^{-1}(\beta(\rho_{0}))\cap{\mathcal{R}}(L)$. Since ${\mathcal{R}}(L)$ is open in ${\mathcal{X}}(S)$, this shows that $(\operatorname{b}^{+})^{-1}(\lambda^{+})$ is locally real-analytic around $\rho_{0}$. Thus it is real-analytic. ∎
@@ 6.2. Prescribed third fundamental form
Given $\rho\in{\mathcal{QF}}_{o}(S)$, a theorem of Labourie [Lab92] states that each end of ${\mathbb{H}}^{3}/\rho(\pi_{1}(S))$ is uniquely foliated by smooth closed subsurfaces of constant curvature equal to $K$, with $K$ contained in the interval (1,0)(-1,0). The third fundamental form of such a subsurface has constant curvature $K^{*}\coloneqq K/(K+1)$. Thus it gives a point in the rescaled Teichmüller space ${\mathcal{T}}^{K^{*}}(S)$, which is the space of isotopy classes of metrics of constant curvature $K^{*}$ on $S$. Focusing on the top end, it follows that, for each $K\in(-1,0)$, one can define the map
(6) $I\hskip-2.84544ptI\hskip-2.84544ptI^{+}_{K}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{T}}^{K^{*}}(S)$
sending a quasifuchsian representation $\rho$ to the third fundamental form of the constant curvature $K$ surface in the top end of ${\mathbb{H}}^{3}/\rho(\pi_{1}(S))$. The third fundamental form $I\hskip-2.84544ptI\hskip-2.84544ptI^{+}_{K}$ is a “smooth measure of bending” and Chen-Schlenker proved that the analogous “smooth version” of Proposition 6.1 holds (see Corollary 1.4 and the proof of Lemma 4.1 in [CS25]). We need the following special case of their result, which is more general.
@@
The map $\operatorname{m}^{-}\times I\hskip-2.84544ptI\hskip-2.84544ptI^{+}_{K}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{T}}(S^{-})\times{\mathcal{T}}^{K^{*}}(S^{+})$ is a homeomorphism.
This is a mixed version of the earlier result of Schlenker that $I\hskip-2.84544ptI\hskip-2.84544ptI^{-}_{K}\times I\hskip-2.84544ptI\hskip-2.84544ptI^{+}_{K}$ is a homeomorphism [Sch06], which was an essential ingredient in [DS24].
@@
When $K\to-1$, the maps $I\hskip-2.84544ptI\hskip-2.84544ptI^{+}_{K}$ converge to $\operatorname{b}^{+}$ in the length spectrum, uniformly on compacts. More precisely, $\ell_{I\hskip-2.84544ptI\hskip-2.84544ptI^{+}_{K}(\rho)}(c)\to i(\lambda^{+}(\rho),c)$ when $K\to-1$, for each closed curve $c$ in $S$ and each $\rho\in{\mathcal{QF}}_{o}(S)$, and the convergence is uniform for $\rho$ in a compact subset of ${\mathcal{QF}}_{o}(S)$.
In order to apply the variant of Finney’s theorem as in the next section, we identify the targets of the maps $I\hskip-2.84544ptI\hskip-2.84544ptI^{+}_{K}$, for $K\in(-1,0)$, and $\operatorname{b}^{+}$, via the earthquake map $E\colon{\mathcal{ML}}(S)\times{\mathcal{T}}(S)\to{\mathcal{T}}(S)\colon(\lambda,m)\mapsto E(\lambda,m)$ as in [DS24, §4.1]. More precisely, we fix a point $m_{0}\in{\mathcal{T}}(S)$ and consider, for each $K\in(-1,0)$, the map
(7) $V_{K}\colon{\mathcal{QF}}_{o}(S)\to{\mathcal{ML}}(S^{+})\colon\rho\mapsto\frac{1}{\sqrt{\left|K^{*}\right|}}E(-,m_{0})^{-1}(\left|K^{*}\right|I\hskip-2.84544ptI\hskip-2.84544ptI^{+}_{K}(\rho)),$
@@ 6.6. Fibres of a limit of homeomorphism
Let $N$ and $N^{\prime}$ be two manifolds of the same dimension. Let $(f_{n}\colon N\to N^{\prime})_{n\in{\mathbb{N}}}$ be a sequence of open embeddings converging continuously (equivalently, compactly) to the continuous map $f_{\infty}\colon N\to N^{\prime}$, i.e. $f_{n}(x_{n})\to f_{\infty}(x_{\infty})$ whenever $x_{n}\to x_{\infty}$. If a fibre of $f_{\infty}$ is compact and is an absolute neighbourhood retract, then it is contractible.
@@
Let $K\coloneqq f_{\infty}^{-1}(\ast)$ be a (nonempty) compact fibre, which is a retract of an open neighbourhood $\Omega$, which we can assume to have compact closure. Let $j\colon K\xhookrightarrow{}\Omega$ be the inclusion and $r\colon\Omega\to K$ a retraction. Assuming that $K$ is not contractible, we will show that $f_{\infty}$ sends a point of $\partial\Omega$ to $\ast$.
Let $U\subset N^{\prime}$ be a contractible neighbourhood of $\ast$. Since $K$ is compact, we have $f_{n}(K)\subset U$ for all $n$ large enough. Fix such a $n\in{\mathbb{N}}$. Thus $f_{n}(K)\subset U$. Since $U$ is contractible, there is a homotopy $H\colon K\times[0,1]\to U$ from $H(-,0)=f_{n}|_{K}\colon K\to U$ to a constant map H(,1)H(-,1). But recall that $f_{n}$ is an open embedding. Thus $H(K\times[0,1])$ cannot be entirely contained in $f_{n}(\Omega)$, for otherwise $r\circ(f_{n}|_{\Omega})^{-1}\circ H\colon K\times[0,1]\to K$ would be a nulhomotopy of $\operatorname{id}_{K}$.
Since $H(K\times[0,1])$ is connected ($H$ is a nulhomotopy, hence every point in its image is connected by a path to the constant image of H(,1)H(-,1)), it must intersect $f_{n}(\partial\Omega)$ (which equals $\partial f_{n}(\Omega)$ since $f_{n}$ is an open embedding), say at a point $f_{n}(x_{U})$ with $x_{U}\in\partial\Omega$. In particular, $f_{n}(x_{U})\in U$.
@@
which is impossible since $\partial\Omega$ does not meet $K=f_{\infty}^{-1}(\ast)$. This concludes the proof that $K$ is contractible. ∎
@@
Since $\operatorname{m}^{-}\times\operatorname{b}^{+}$ is the limit of the homeomorphisms $\operatorname{m}^{-}\times V_{K}$ and the convergence is uniform on compacts, $F$ is contractible by Theorem 6.7.
Since $F$ is a compact real-analytic space, it has a fundamental class modulo two (by Borel-Haefliger and Sullivan, see Theorem 5.3), i.e. a nonzero homology class $\alpha\in H_{\dim(F)}(F;{\mathbb{Z}}/2)$. But $F$ is contractible, hence its homology is trivial in all degrees except 0. Since $\alpha\neq 0$, we must have $\dim(F)=0$. Therefore, $F$ is a contractible compact real-analytic space of dimension 0, that is, a point. This proves that $\operatorname{m}^{-}\times\operatorname{b}^{+}$ is injective. By invariance of domains, it is an open embedding. This concludes, since a proper open embedding into a connected space is a homeomorphism. ∎
@@ 6.8. Conclusion
is proper (Proposition 4.1), continuous (on the first factor, by the simultaneous uniformization theorem, see Section 2.10 ; on the second factor, by Proposition 3.4), and its domain and target are manifolds of the same dimension (By [Ker92, Theorem 4.7], ${\mathcal{ML}}^{\perp e^{-}}_{<\pi}(S^{+})$ is a manifold of the same dimension as ${\mathcal{T}}(S^{+})$), it suffices to show that it is injective. With the notation of Section 1.1, it suffices to show that $\operatorname{b}^{+}_{\infty}$ is injective. But it is the limit of the open embeddings $\operatorname{b}^{+}_{t}$, when $t\to\infty$, and the convergence is continuous since $\operatorname{B}^{+}$ is continuous. By following the proof of Proposition 6.1 and using Theorem 6.7, we obtain that $\operatorname{b}^{+}_{\infty}$ is injective, hence $\operatorname{b}^{+}$ is injective too. ∎